# No Representation, No Trust: Connecting Representation, Collapse, and Trust Issues in PPO

### v1

@@ Abstract
Reinforcement learning (RL) is inherently rife with non-stationarity since the states and rewards the agent observes during training depend on its changing policy. Therefore, networks in deep RL must be capable of adapting to new observations and fitting new targets. However, previous works have observed that networks in off-policy deep value-based methods exhibit a decrease in representation rank, often correlated with an inability to continue learning or a collapse in performance. Although this phenomenon has generally been attributed to neural network learning under non-stationarity, it has been overlooked in on-policy policy optimization methods which are often thought capable of training indefinitely. In this work, we empirically study representation dynamics in Proximal Policy Optimization (PPO) on the Atari and MuJoCo environments, revealing that PPO agents are also affected by feature rank deterioration and loss of plasticity. We show that this is aggravated with stronger non-stationarity, ultimately driving the actor’s performance to collapse, regardless of the performance of the critic. We draw connections between representation collapse, performance collapse, and trust region issues in PPO, and present Proximal Feature Optimization (PFO), a novel auxiliary loss, that along with other interventions shows that regularizing the representation dynamics improves the performance of PPO agents. Code and run histories are available at https://github.com/CLAIRE-Labo/no-representation-no-trust.
@@ 1 Introduction
Reinforcement learning (RL) agents are inherently subject to non-stationarity as the states and rewards they observe change during learning. Therefore, neural networks in deep RL must be capable of adapting to new inputs and fitting new targets. However, previous works have observed that value networks in off-policy value-based algorithms exhibit a decrease in the rank of their representations, termed feature rank, and a decrease in their ability to regress to arbitrary targets, called plasticity (Kumar et al., 2021; Lyle et al., 2022). Consequently, this decrease in representation fitness was often correlated with the value network’s inability to continue learning and adapting to new tasks, especially in sparse-reward scenarios. Although this phenomenon is more generally attributed to neural networks trained under non-stationarity (Lyle et al., 2023), it has been overlooked in on-policy policy optimization methods that are often thought capable of training indefinitely. In particular, Proximal Policy Optimization (PPO) (Schulman et al., 2017), one of the most popular policy optimization methods, introduces additional non-stationarity by making several minibatch updates over non-stationary data and by optimizing a surrogate loss depending on a moving policy. This raises the question of how much PPO agents are impacted by the same representation degradation attributed to non-stationarity. Igl et al. (2021) have shown that non-stationarity affects the generalization of PPO agents (learning speed when training episodes are very different otherwise performance at test time on novel episodes) but does not prevent training, and no connection was made with the feature rank and plasticity measures used in the recent value-based works.
In this work, we conduct an empirical study of representation dynamics in PPO, highlighting the detrimental effects of non-stationarity on policy optimization. Our contributions are the following:
@@
    We provide the first study of feature rank and plasticity in policy optimization, revealing that PPO agents in the Arcade Learning Environment (Bellemare et al., 2013) and MuJoCo (Todorov et al., 2012) environments are subject to representation collapse.
@@
    We draw connections between representation collapse, performance collapse, and trust region issues in PPO, showing that PPO’s clipping becomes ineffective under poor representations and fails to prevent performance collapse, which is irrecoverable due to loss of plasticity.
@@
    We propose Proximal Feature Optimization (PFO), a new regularization on the representation of the policy that regularizes the change in pre-activations. By addressing the representation issues, PFO can mitigate performance collapse and improve the agent’s performance.
@@
    We open source our code and run histories, providing a comprehensive and reproducible codebase for studying representation dynamics in policy optimization and a large database of run histories with extensive logging for further investigation on this topic.
@@ 2 Background
We consider on-policy deep actor-critic agents which train a policy network $\pi(\cdot;{\bm{\theta}})$, called actor and denoted $\pi_{\bm{\theta}}$, and a value network $\hat{v}(\cdot;{\bf w})$, called critic that approximates the return of $\pi_{\bm{\theta}}$ at every state (expected sum of rewards starting from a state). At every training stage, the agent collects a batch of samples, called rollout, with its current policy $\pi_{\bm{\theta}}$, and both networks are trained with gradient descent on this data. The critic is trained to minimize the Euclidean distance to an estimator of the returns (e.g., $G_{t}$). We use $\lambda$-returns computed with the Generalized Advantage Estimator (GAE) (Schulman et al., 2015b). The actor is trained with the Proximal Policy Optimization (PPO) (Schulman et al., 2017).
@@
PPO-Clip, the most popular variant of PPO algorithms proposed by Schulman et al. (2017), optimizes the actor by maximizing the surrogate objective in Equation 1 at each rollout.
@@
The objective is defined for some small hyperparameter $\epsilon$, where $\pi_{\text{old}}$ is the last $\pi_{\bm{\theta}}$ of the previous optimization stage and is used to collect the training batch, and $\Psi_{t}$ is an estimator of the advantage of $\pi_{\text{old}}$ (e.g., ${\Psi_{t}=G_{t}-\hat{v}(S_{t};{\bf w}_{\text{old}})}$). We use the GAE in our experiments. The objective is optimized with minibatch gradient steps over multiple epochs on the rollout data (batch). We refer to PPO-Clip as PPO and provide a high-level pseudocode in Algorithm 1.
Intuitively, PPO aims to maximize the policy advantage $\mathbb{E}_{\pi_{\text{old}}\!\!}\left[\sum_{t=0}^{t_{\max}-1}\frac{\pi_{\bm{\theta}}(A_{t}|S_{t})}{\pi_{\text{old}}(A_{t}|S_{t})}\Psi_{t}\right]$ defined by Kakade & Langford (2002), which participates in a lower bound to the improvement of $\pi_{\bm{\theta}}$ given that it is close to $\pi_{\text{old}}$ (Schulman et al., 2015a, see Theorem 1). In this regard, a gradient step on $L_{\pi_{\text{old}}}^{CLIP}(\theta)$ would increase (decrease) the probability of actions at states yielding positive (negative) advantage until the ratio between the policies for those actions reaches $1+\epsilon$ ($1-\epsilon)$ at which point the gradient at those samples becomes null. This is a heuristic to ensure a trust region that keeps policies close to each other, resulting in policy improvement.
@@
The actor and the critic networks are both subject to non-stationarity in deep RL. As the agent improves, it visits different states, shifting the distribution of states which makes the networks’ input distribution non-stationary. This also holds for the targets to fit the critic, which change as the policy returns change. This form of non-stationarity has been shown by previous work to hinder the ability of several deep RL agents (e.g., DQN (Mnih et al., 2015), SAC (Haarnoja et al., 2018)) to continue learning (Lyle et al., 2022; Nikishin et al., 2022). PPO introduces additional non-stationarity to the actor compared to policy gradient methods (such as vanilla policy gradient (Sutton et al., 1999), A2C Mnih et al. (2016)) by making its objective non-stationary as it depends on the previous policy. In addition, the practical benefit of the PPO objective is that it can be optimized by performing multiple epochs of mini-batch gradient descent on every new collected batch; however, this makes the networks more likely to be impacted by previous training rollouts as new rollouts are collected. Increasing the number of epochs can cause the agent to “overfit” more to previous experience, making the impact of non-stationarity stronger.
@@
As done in most works studying feature dynamics in deep RL (Lyle et al., 2022; Kumar et al., 2021), we refer to the activations of the last hidden layer of a network (the penultimate layer) as the features or representation learned by the network. On a batch of $N$ samples, this gives a matrix of dimension $N\times D$ denoted by $\Phi$, where $D<N$ is the width of the penultimate layer. Several measures of the rank of this matrix have been used to quantify the “quality” of the representation (Kumar et al., 2021; Gulcehre et al., 2022; Lyle et al., 2022; Andriushchenko et al., 2023). Their absolute values differ significantly, but their dynamics are often correlated. We track all of the different rank metrics in our experiments, compare them in Appendix E.1 and E, and use the approximate rank in our main figures for its connection to principal component analysis (PCA). Given a threshold $\delta\in\mathbb{R}$ and the singular values of the feature matrix $\langle\sigma_{i}(\Phi),\dots,\sigma_{D}(\Phi)\rangle$ in decreasing order, the approximate rank of $\Phi$ is $\min_{k}\left\{\frac{\sum_{i=1}^{k}\sigma_{i}^{2}(\Phi)}{\sum_{j=1}^{D}\sigma_{j}^{2}(\Phi)}>1-\delta\right\}$ which corresponds to the smallest dimension of the subspace recovering $(1-\delta)\%$ of the variance of $\Phi$. We use $\delta=0.01$ i.e. the reconstruction recovers $99\%$ of the variance as done by Andriushchenko et al. (2023); Yang et al. (2020). We also refer to this metric as feature rank with reference to the rank of the feature matrix when there is no ambiguity.
Plasticity loss
Plasticity is computed on checkpoints of a network undergoing some training to measure the evolution of its ability to fit targets. Given a fixed target and a fixed optimization budget (number of gradient steps), a checkpoint’s plasticity loss is the loss from fitting the checkpoint to the target at the end of the optimization budget. Usually, the plasticity of a deep RL agent throughout training is measured by its ability to fit the outputs of a model initialized from the same distribution as the agent, on a fixed rollout collected by this target random model (Lyle et al., 2022; Nikishin et al., 2023). The data would in expectation be from the same distribution as the agent’s initial checkpoint. To fit the critic we use an $L^{2}$ loss on the outputs of the models. To fit the actor we use a KL divergence between the target and the checkpoint (forward KL).
@@ 3 Deteriorating representations, collapse, and loss of trust
It is well-known that non-stationarity in deep RL can be a factor causing issues in representation learning. However, most of the observations have been made in value-based methods showing that value networks are prone to rank collapse, harming their expressivity, and in turn, the performance of the agent (Lyle et al., 2022; Kumar et al., 2022); Igl et al. (2021) studied non-stationarity in PPO but only showed that it harms its generalization, with no evidence of rank deterioration or performance collapse. Our motivation is to reuse the tools that showed that value-based methods are prone to representation collapse but in policy optimization methods for the first time. We focus on PPO for its popularity and extra non-stationarity through multi-epoch optimization.
@@
We begin our experiments by training PPO agents on the Arcade Learning Environment (ALE)(Bellemare et al., 2013) for pixel-based observations with discrete actions and on MuJoCo (Todorov et al., 2012) for continuous observations with continuous actions. To keep our experiments tractable, we choose the Atari-5 subset recommended by Aitchison et al. (2023) and add Gravitar to include at least one sparse-reward hard-exploration game from the taxonomy presented by Bellemare et al. (2016). For MuJoCo, we train on Ants, Half-Cheetahs, Humanoids, and Hoppers, which have varying complexity and observation and output sizes. We use the same model architectures and hyperparameters as popular implementations of PPO on ALE and MuJoCo (Raffin et al., 2021; Huang et al., 2022b); these are also the architectures and hyperparameters used by Schulman et al. (2017) in the original implementation of PPO, they do not include normalization layers. The ALE model uses ReLU activations (Nair & Hinton, 2010) and the MuJoCo one tanh; we also experiment with ReLU on MuJoCo. We use separate actor and critic models for both environments unless specified in Section 4. Details on the performance metrics and tables of all environment parameters, model architectures, and algorithm hyperparameters are presented in Appendix B. Observing that the previous findings on the feature dynamics of value-based approaches (Gulcehre et al., 2022; Lyle et al., 2022) apply to the critic of PPO as well since the loss function is the same, we focus on studying the feature dynamics of the actor unless stated otherwise in the text or figures.
@@
We vary the number of epochs as a way to control the effect of non-stationarity, which gives the agent a more significant number of optimization steps per rollout while not changing the optimal target it can reach due to clipping, as opposed to changing the value of $\epsilon$ in the trust region for example. We keep the learning rate constant throughout training and use the same value for all the epoch configurations. To understand the feature dynamics, we measure different metrics that are proposed in the literature, including feature rank, number of dead neurons (Gulcehre et al., 2022), plasticity loss (Lyle et al., 2022), and penultimate layer pre-activation norm. Previous work has monitored feature norm values as well (Lyle et al., 2024); however, in our case, we found that as the neurons in the policy network die, the feature norm might be stable while the pre-activation norm blows up. All the metrics are computed on on-policy rollouts except for the plasticity loss.
@@
We run five seeds per hyperparameter configuration and report mean curves with min/max shaded regions unless specified otherwise. All curves, except for plasticity loss, are smoothed using an exponentially weighted moving average of coefficient $0.05$. We seek to answer the following questions:
  • Q1. How do a PPO agent’s representation metrics, such as the feature rank and the plasticity loss, evolve during training? Are they subject to the same decline observed by Kumar et al. (2021); Lyle et al. (2022) in value-based methods? Does it affect performance?
  • Q2. How does increasing the number of epochs per rollout to vary non-stationarity affect a PPO agent’s representation? Does it degrade as observed in DQN and SAC agents when increasing the replay ratio (Nikishin et al., 2022; Kumar et al., 2022)?

### v3

@@ Abstract
Reinforcement learning (RL) is inherently rife with non-stationarity since the states and rewards the agent observes during training depend on its changing policy. Therefore, networks in deep RL must be capable of adapting to new observations and fitting new targets. However, previous works have observed that networks trained under non-stationarity exhibit an inability to continue learning, termed loss of plasticity, and eventually a collapse in performance. For off-policy deep value-based RL methods, this phenomenon has been correlated with a decrease in representation rank and the ability to fit random targets, termed capacity loss. Although this correlation has generally been attributed to neural network learning under non-stationarity, the connection to representation dynamics has not been carefully studied in on-policy policy optimization methods. In this work, we empirically study representation dynamics in Proximal Policy Optimization (PPO) on the Atari and MuJoCo environments, revealing that PPO agents are also affected by feature rank deterioration and capacity loss. We show that this is aggravated by stronger non-stationarity, ultimately driving the actor’s performance to collapse, regardless of the performance of the critic. We ask why the trust region, specific to methods like PPO, cannot alleviate or prevent the collapse and find a connection between representation collapse and the degradation of the trust region, one exacerbating the other. Finally, we present Proximal Feature Optimization (PFO), a novel auxiliary loss that, along with other interventions, shows that regularizing the representation dynamics mitigates the performance collapse of PPO agents. Code and run histories are available at https://github.com/CLAIRE-Labo/no-representation-no-trust.
@@ 1 Introduction
Reinforcement learning (RL) agents are inherently subject to non-stationarity as the states and rewards they observe change during learning. Therefore, neural networks in deep RL must be capable of adapting to new inputs and fitting new targets. However, previous works have observed that networks trained under non-stationarity exhibit an inability to continue learning, termed loss of plasticity, and a collapse in performance (Dohare et al., 2021; Abbas et al., 2023; Kumar et al., 2023; Dohare et al., 2023a, b). Kumar et al. (2021); Lyle et al. (2022) connect this phenomenon to representation dynamics and show that value networks in off-policy value-based RL algorithms exhibit a decrease in the rank of their representations, termed feature rank collapse, and a decrease in their ability to regress to arbitrary targets, called capacity loss. Although this deterioration in representation is more generally attributed to neural networks trained under non-stationarity (Lyle et al., 2023), the connection to representation dynamics has been overlooked in on-policy policy optimization methods. In particular, Proximal Policy Optimization (PPO) (Schulman et al., 2017), one of the most popular policy optimization methods, makes several minibatch updates over non-stationary data, unlike vanilla policy gradient methods, and optimizes a surrogate loss that depends on a moving old policy. This raises the question of how much PPO agents are impacted by the same representation degradation attributed to non-stationarity.  Dohare et al. (2021, 2023a, 2023b) have shown that PPO agents lose plasticity throughout training but have only diagnosed it as a collapse in performance or as an Adam optimization issue. Igl et al. (2021) have shown that non-stationarity affects the generalization of PPO agents (learning speed when training episodes are very different otherwise performance at test time on novel episodes) but not necessarily training, and no connection was made with the feature rank and capacity measures used in the recent value-based works. One crucial outstanding question is why the trust region embedded in methods like PPO is unable to prevent the deterioration in policy by constraining its update. To address these gaps, we present the following contributions:
@@
    We provide the first study of feature rank and capacity loss in on-policy policy optimization, revealing that PPO agents in the Arcade Learning Environment (Bellemare et al., 2013) and MuJoCo (Todorov et al., 2012) environments are subject to representation collapse.
@@
    We draw connections between representation collapse, performance collapse, and trust region issues in PPO, showing that PPO’s clipping becomes ineffective under poor representations and fails to prevent performance collapse, which is irrecoverable due to loss of capacity. We further isolate the breakdown of the trust region in a theoretical setting.
@@
    We propose Proximal Feature Optimization (PFO), a new regularization on the representation of the policy that regularizes the change in pre-activations. PFO strengthens our analysis by addressing the representation issues and mitigating performance collapse.
@@
    We open source our code providing a comprehensive and reproducible codebase for studying representation dynamics in policy optimization and a large database of run histories with extensive logging for further investigation on this topic.
@@ 2 Background
We consider on-policy deep actor-critic agents which train a policy (or actor) network $\pi(\cdot;{\bm{\theta}})$ also denoted $\pi_{\bm{\theta}}$, and a value (or critic) network $\hat{v}(\cdot;{\bf w})$ that approximates the return of $\pi_{\bm{\theta}}$ at every state. At every training stage, the agent collects a batch of samples, called rollout, with its current policy $\pi_{\bm{\theta}}$, and both networks are trained with gradient descent on this data. The critic is trained to minimize the Euclidean distance to an estimator of the returns (e.g., $G_{t}$). We use $\lambda$-returns computed with the Generalized Advantage Estimator (GAE) (Schulman et al., 2015b). The actor is trained with the Proximal Policy Optimization (PPO) (Schulman et al., 2017).
@@
PPO-Clip, the most popular variant of PPO algorithms (Schulman et al., 2017), optimizes the actor by repeatedly maximizing the objective in Equation 1 at each rollout.
@@
The objective is defined for some small hyperparameter $\epsilon$; $\pi_{\text{old}}$ is the last $\pi_{\bm{\theta}}$ of the previous optimization phase, used to collect a training batch (rollout) after each optimization phase; $\Psi_{t}$ is an estimator of the advantage of $\pi_{\text{old}}$ (e.g., ${\Psi_{t}=G_{t}-\hat{v}(S_{t};{\bf w}_{\text{old}})}$); we use the GAE in our experiments. An optimization phase consists of maximizing the objective with minibatch gradient steps over multiple epochs on the training batch. We refer to PPO-Clip as PPO and provide a pseudocode in Algorithm 1.
Intuitively, PPO aims to maximize the policy advantage $\mathbb{E}_{\pi_{\text{old}}\!\!}\left[\sum_{t=0}^{t_{\max}-1}\frac{\pi_{\bm{\theta}}(A_{t}|S_{t})}{\pi_{\text{old}}(A_{t}|S_{t})}\Psi_{t}\right]$ defined by Kakade & Langford (2002), which participates in a lower bound to the improvement of $\pi_{\bm{\theta}}$ given that it is close to $\pi_{\text{old}}$ (Schulman et al., 2015a, see Theorem 1). In this regard, a gradient step on $L_{\pi_{\text{old}}}^{CLIP}(\theta)$ would increase (resp. decrease) the probability of actions at states yielding positive (resp. negative) advantage until the ratio between the policies for those actions reaches $1+\epsilon$ (resp. $1-\epsilon)$ at which point the gradient at those samples becomes null. This is a heuristic to ensure a trust region that keeps policies close to each other, resulting in policy improvement.
@@
The actor and the critic networks are both subject to non-stationarity in deep RL. As the agent improves, it visits different states, shifting the distribution of states which makes the networks’ input distribution non-stationary. This also holds for the targets to fit the critic, which change as the returns of the policy change. Unlike vanilla policy gradient (Sutton et al., 1999), and A2C (Mnih et al., 2016), PPO’s objective is optimized by performing multiple epochs of minibatch gradient descent on the current collected batch, potentially making the networks more likely to be impacted by previous training rollouts. In this sense, increasing the number of epochs in PPO can cause the agent to “overfit” more to previous experience.
@@
As done in most works studying feature dynamics in deep RL (Lyle et al., 2022; Kumar et al., 2021), we refer to the activations of the last hidden layer of a network (the penultimate layer) as the features or representation learned by the network. On a batch of $N$ samples, this gives a matrix of dimensions $N\times D$ denoted by $\Phi$, where $D<N$ is the width of the penultimate layer. Several measures of the rank of this matrix have been used to quantify the “quality” of the representation (Kumar et al., 2021; Gulcehre et al., 2022; Lyle et al., 2022; Andriushchenko et al., 2023). Their absolute values differ significantly, but their dynamics are often correlated. We track all of the different rank metrics in our experiments, compare them in Appendix E, and use the approximate rank in our main figures for its connection to principal component analysis (PCA). Given a threshold $\delta\in\mathbb{R}$ and the singular values $\langle\sigma_{i}(\Phi),\dots,\sigma_{D}(\Phi)\rangle$ of $\Phi$ in decreasing order, the approximate rank of $\Phi$ is $\min_{k}\left\{\frac{\sum_{i=1}^{k}\sigma_{i}^{2}(\Phi)}{\sum_{j=1}^{D}\sigma_{j}^{2}(\Phi)}>1-\delta\right\}$ which corresponds to the smallest dimension of the subspace recovering $(1-\delta)\%$ of the variance of $\Phi$. We use $\delta=0.01$ i.e. the reconstruction recovers $99\%$ of the variance as done by Andriushchenko et al. (2023); Yang et al. (2020). We refer to this metric as feature rank with reference to the rank of the feature matrix when there is no ambiguity.
Capacity loss
Target-fitting capacity (Lyle et al., 2022) is computed on checkpoints of a network undergoing some training to measure the evolution of its ability to fit some chosen target independent from its training. It is a concrete metric to evaluate plasticity. Given a fixed target (distribution over inputs and outputs) and a fixed optimization budget, a checkpoint’s capacity loss is the loss from fitting the checkpoint to the target at the end of the optimization budget. Usually, the capacity of a deep RL agent is measured by its ability to fit the outputs of a model initialized randomly from the same distribution as the agent on a fixed rollout collected by this target random model (Lyle et al., 2022; Nikishin et al., 2023). We follow this practice. The data would in expectation be from the same distribution as the agent’s initial checkpoint. To fit the critic, we use an $L^{2}$ loss on the outputs of the models. To fit the actor, we use a KL divergence between the target and the checkpoint (forward KL).
@@ 3 Deteriorating representations, collapse, and loss of trust
It is well-known that non-stationarity in deep RL can be a factor causing issues in representation learning. However, most of the observations have been made in value-based methods showing that value networks are prone to rank collapse, harming their expressivity, and in turn, the performance of the agent (Lyle et al., 2022; Kumar et al., 2022). Non-stationarity has been shown to impact PPO’s generalization Igl et al. (2021) and performance in the long run or in a continual learning setting (Dohare et al., 2021, 2023a, 2023b), but no evidence of representation deterioration was shown. Our motivation is to reuse the tools that showed that value-based methods are prone to representation collapse but in policy optimization methods for the first time. We focus on PPO for its popularity and its non-stationarity which is impacted and can be controlled by multi-epoch optimization. Furthermore, a crucial question for PPO, compared to most value-based alternatives, is how the regularization implicit in PPO through its trust region interacts with representation and performance collapse. Intuitively it should prevent rapid degradation of the policy.
@@
We begin our experiments by training PPO agents on the Arcade Learning Environment (ALE)(Bellemare et al., 2013) for pixel-based observations with discrete actions and on MuJoCo (Todorov et al., 2012) for continuous observations with continuous actions. To keep our experiments tractable, we choose the Atari-5 subset recommended by Aitchison et al. (2023) and add Gravitar to include at least one sparse-reward hard-exploration game from the taxonomy presented by Bellemare et al. (2016). For MuJoCo, we train on Ants, Half-Cheetahs, Humanoids, and Hoppers, which have varying complexity and observation and output sizes. We use the same model architectures and hyperparameters as popular implementations of PPO on ALE and MuJoCo (Raffin et al., 2021; Huang et al., 2022b); these are also the architectures and hyperparameters used by Schulman et al. (2017) in the original implementation of PPO; they do not include normalization layers. For MuJoCo we further adopt a parameterization of the output action distribution using a TanhNormal [^3] We also provide evidence of collapse with the Gaussian distribution parameterization in the appendix. with both its mean and variance depending on the state representation as done by Haarnoja et al. (2018); Andrychowicz et al. (2021). As we study the connection between performance and representation dynamics this is a more natural choice than using the commonly implemented state-independent variance which would be independent of representation dynamics. The ALE models use ReLU activations (Nair & Hinton, 2010) and the MuJoCo ones tanh; we also experiment with ReLU on MuJoCo. We use separate actor and critic models for both environments unless specified in Section 4. Details on the performance metrics and tables of all environment parameters, model architectures, and algorithm hyperparameters are presented in Appendix B. Observing that the previous findings on the feature dynamics of value-based approaches (Gulcehre et al., 2022; Lyle et al., 2022) apply to the critic of PPO as well since the loss function is the same, we focus on studying the feature dynamics of the actor unless stated otherwise in the text or figures.
@@
We vary the number of epochs as a tool to control the effects of non-stationarity, which gives the agent a more significant number of optimization steps per rollout while not changing the optimal target it can reach due to clipping, as opposed to changing the value of $\epsilon$ in the trust region for example.444 We show results with varying $\epsilon$ in Figure 34 of Appendix D. We keep the learning rate constant throughout training and use the same learning rate for all the epoch configurations. [^5] Although the environments we use are single-task environments to ablate additional MDP non-stationarity, they are complex enough for the agents to keep improving when trained for longer than common benchmark limits without annealing the learning rate. To understand the feature dynamics, we measure different metrics that are proposed in the literature, including feature rank, number of dead neurons (Gulcehre et al., 2022), capacity loss (Lyle et al., 2022), and penultimate layer pre-activation norm. Previous work has monitored feature norm values as well (Abbas et al., 2023; Lyle et al., 2024); however, in our case, we found that as the neurons in the policy network die, the feature norm might be stable while the pre-activation norm blows up. All the metrics are computed on on-policy rollouts except for the capacity loss.
@@
We run five seeds per hyperparameter configuration and report mean curves with min/max shaded regions unless specified otherwise. All curves, except for capacity loss, are smoothed using an exponentially weighted moving average with a coefficient of $0.05$.